OpenStudy (anonymous):

What is the equation of the line which includes points (-3, 2) and (1, 6)? A. y = x - 5 B. y = x + 5 C. D.

OpenStudy (anonymous):

Without doing the whole question out, you can just plug in both sets of points to see what works: A) 2 = -3 - 5 2 = - 8 Therefore A does not work. B) 2 = -3 + 5 2 = 2 6 = 1 + 5 6 = 6 B works with both points, therefore B. (y = x +5) is the answer. Hope this helped!